Earthquakes generate two circular waves, p-waves and s-waves, tht travel outwards. the p-waves travels faster, but generally the s-waves do the most damage. suppose the p-waves have a velocity of 3.6 miles per second and the s-waves have a velocity of 1.8 miles per second. How long after the earthquake will p-waves and s-waves be 80 miles apart?
Answer by ptaylor(2198)


distance(d)=rate(r) times time(t) or d=rt; t=d/r
Now we know that when the distance the p-wave travels minus the distance the s-wave travels equals 80 mi, we will have the time we are looking for. So
rt of p wave=3.6t
rt of s wave=1.8t
3.6t-1.8t=80
1.8t=80 divide both sides by 1.8
t=44.4444 sec
ck
3.6(44.44444)-1.8(44.44444)=80
~80=80

Hope this helps----ptaylor